Circular No. 7983

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 T1 (LINEAR) M. Blythe and S. Partridge, Lincoln Laboratory, Massachusetts Insti-

tute of Technology, report the LINEAR discovery of a comet showing a coma and a possible tail in p.a. ∼ 260^{o} (discovery observation below). Following posting on the NEO Confirmation Page, other observers confirmed the cometary appearance, including P. Birtwhistle, Great Shefford, England (Oct. 4.11 UT, m_{1} = 15.4, faint 1′ tail in p.a. 231^{o}), and J. McGaha, Tucson, AZ (Oct. 4.33, nuclear condensation of mag ∼ 15.3 and diameter ∼ 6′′, with a tail 45′′ long and ∼ 12′′ wide in p.a. 235^{o} curving toward the north).

2002 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 3.25540  1^{h}43^{m}00.^{s}66  +2^{o}56′58′′.5  16.2

The available astrometry, the following preliminary elliptical elements, and an ephemeris appear on MPEC 2002-T15.

T = 2002 Oct. 8.030 TT  ω = 1.^{o}256
e = 0.65413  Ω = 15.485 }2000.0
q = 1.19004 AU  i = 20.528
a = 3.44071 AU n^{o} = 0.154430 P = 6.38 years

SUPERNOVAE 2002gb AND 2002gc Two apparent supernova discoveries have been reported. Further to

IAUC 7979, W. M. Wood-Vasey, G. Aldering, and P. Nugent report their discovery of SN 2002gb on unfiltered Palomar NEAT images. T. Boles, Coddenham, England, reports his discovery of SN 2002gc in UGC 1394 on an unfiltered CCD image taken with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol.

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002gb Aug. 28.26 22^{h}43^{m}21.^{s}31 − 0^{o}06′53′′.0 18.9 1′′.4 S
2002gc Oct. 3.88 1 55 46.56 +46 48 19.4 17.4 44′′.3 E, 11′′.9 N

Additional unfiltered magnitudes reported by the respective observers: SN 2002gb, 2001 Oct. 5 UT, [20.5; 2002 Aug. 16, [20.5; Sept. 9, 19.4; Oct. 3, 19.5. SN 2002gc, 2001 Sept. 3, [19.0; 2002 July 15, not present; Aug. 12, not present; Oct. 4.845, 17.3. Boles adds that SN 2002gc is not present on second Palomar Sky Survey red and blue plates.

2002 October 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green